Exercise 13-06

The payroll of Skysong Company for September 2019 is as follows.

Total payroll was $450,000, of which $118,000 is exempt from Social Security tax because it represented amounts paid in excess of $128,400 to certain employees. The amount paid to employees in excess of $7,000 was $439,000. Income taxes in the amount of $77,200 were withheld, as was $9,600 in union dues. The state unemployment tax is 3.5%, but Skysong Company is allowed a credit of 2.3% by the state for its unemployment experience. Also, assume that the current FICA tax is 7.65% on an employee’s wages to $128,400 and 1.45% in excess of $128,400. No employee for Skysong makes more than $125,000. The federal unemployment tax rate is 0.8% after state credit.

Prepare the necessary journal entries if the wages and salaries paid and the employer payroll taxes are recorded separately. (Round answers to 0 decimal places, e.g. 5,275. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts .Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Account Titles and Explanation

Debit

Credit

enter an account title to record wages and salaries paid

enter a debit amount

enter a credit amount

enter an account title to record wages and salaries paid

enter a debit amount

enter a credit amount

enter an account title to record wages and salaries paid

enter a debit amount

enter a credit amount

enter an account title to record wages and salaries paid

enter a debit amount

enter a credit amount

enter an account title to record wages and salaries paid

enter a debit amount

enter a credit amount

(To record wages and salaries paid)

enter an account title to record employer payroll taxes

enter a debit amount

enter a credit amount

enter an account title to record employer payroll taxes

enter a debit amount

enter a credit amount

enter an account title to record employer payroll taxes

enter a debit amount

enter a credit amount

enter an account title to record employer payroll taxes

enter a debit amount

enter a credit amount

(To record employer payroll taxes)
Answer #1
Account Titles and Explanation Debit Credit
Salaries and Wages Expense 450000
    Withholding Income Taxes Payable 77200
     FICA Taxes Payable 27109 =(450000-118000)*7.65%+(118000*1.45%)
     Union Dues Payable 9600
     Cash 336091
(To record wages and salaries paid)
Payroll Tax Expense 27329
     FICA Taxes Payable 27109
     Federal unemployment(FUTA) Taxes Payable 88 =(450000-439000)*0.8%
     State unemployment (SUTA) Taxes Payable 132 =(450000-439000)*(3.5%-2.3%)
(To record employer payroll taxes)